OpenStudy (anonymous):

how to solve 30/n-7n+12=-25

OpenStudy (igreen):

Is this the equation? \(\dfrac{30}{n} - 7n + 12 = -25\)

OpenStudy (igreen):

@kateramos

OpenStudy (anonymous):

u have to isolate the variable to get n=?

OpenStudy (anonymous):

Yes, I need to isolate the variable. In trying to cancel out the n in the demoninator, I end up with 7n being squared. I don't want this to happen. What is your suggestion?

OpenStudy (anonymous):

hol on i will shpw u

OpenStudy (swamijatin2):

30/n-7n+12=-25 can u take LCM of 30/n-7n

OpenStudy (anonymous):